At a baseball game, a vender sold a combined total of 195 sodas and hot dogs. The number of hot dogs sold was 45 less than the number of sodas sold. Find the number of sodas sold and the number of hot dogs sold.
Answer by ikleyn(52781)

.

    H + (H + 45) = 195

       2H        = 195 - 45 = 150

        H                   = 150/2 = 75.


ANSWER.  75 hot dogs.

Solved.